OpenStudy (anonymous):

Solve the quadratic equation by factorisation 3x^2 - 7x + 2 = 0 Having trouble with this one!

OpenStudy (anonymous):

(3x-1)(x-2)

OpenStudy (anonymous):

f(x) = 3x^2 - 7x + 2 f(2) = 0 (3x^2-7x+2)/(x-2) = 3x-1 therefore 3x^2-7x+2 = (x-2)(3x-1) x = 2 or x=1/3

OpenStudy (anonymous):

The answer is actually to solve , by doing as SaindDavid did is just factorization.

OpenStudy (anonymous):

should we provide the actual answers all the time, or rather help these guys out on figuring it out?

OpenStudy (anonymous):

so the x terms have to multiply to get 3x^2. Luckily the only factors of 3 are 3 and one, so we have (3x + a)(x + b). We know ab has to equal 2, and 3b + a has to equal -7, because we know how it multiplies out. So a is -2 and b is -1. So we have the answer from above: (3x - 1)(x - 2) = 0 x = 2 or x = 1/3

OpenStudy (anonymous):

i use the quadratic formula . and got the same answer as saintdavid :) \[-b \pm \sqrt{b^{2}-4ac}/2a\]

OpenStudy (anonymous):

@SaintDavid91 and @Mathlearner22, you say this but I believe most people actually show it step by step so the one asking the question understands how it has been done. Instead on just "commenting" without actually helping.....

OpenStudy (anonymous):

sorry :)